# Reshaping Neural Representation via Associative, Presynaptic Short-Term Plasticity

### v1

@@
[…]
Genki Shimizu, Taro Toyoizumi
Short-term synaptic plasticity (STP) is traditionally viewed as a purely presynaptic mechanism that filters incoming spike trains without detecting postsynaptic activity. Recent experiments, however, have revealed an associative form of STP in which presynaptic release probability is modified in concert with long-term potentiation, suggesting a richer computational role for presynaptic plasticity. Here we develop a normative theory of such associative STP using an information-theoretic framework. Extending Fisher-information-based learning rules to synapses with Tsodyks–Markram dynamics, we derive analytic update rules for both baseline synaptic strength and release probability that maximize encoded stimulus information under resource constraints. The resulting learning rules decompose into a conventional postsynaptic term that tracks local firing and a distinct presynaptic term. The latter term exhibits a novel phase-advanced structure relative to firing rate and selectively detects stimulus onset; critically, it is within this presynaptic component that the plasticity of baseline strength and release probability differ. When stimulus drive varies on timescales slower than the EPSP time constant, this onset sensitivity biases optimal connectivity toward anti-causal associations, strengthening synapses from neurons activated later in a stimulus to those activated earlier. In recurrent circuits, networks optimized under these rules develop ramp-like representations during sustained stimulation and support reverse replay sequences when external drive is removed. Linear-response analysis further shows that STP endows presynaptic drive with frequency-dependent phase selectivity and that constraints on total release probability systematically modulate the strength of temporal asymmetry. Together, these results provide a principled account of the computational role of associative STP and highlight presynaptic plasticity of release probability as a key substrate for rapidly reconfigurable temporal coding.
@@ 1 Introduction
Learning through experience is mediated by synaptic plasticity, the activity-dependent modification of connection strengths between neurons. Classical long-term potentiation (LTP) has been extensively studied as the primary mechanism for associative learning, where correlated pre- and postsynaptic activity leads to persistent strengthening of synapses through increased postsynaptic receptor expression [bliss1973Longlasting, bi1998Synaptic]. This Hebbian plasticity enables the formation of neural assemblies that encode relationships between concepts. However, LTP typically requires tens of minutes to hours for full expression and once expressed, can persist from hours to days or longer, which cannot fully account for flexible behavioral adaptation occurring on faster timescales.
@@
To address this question, we adopt an information-theoretic framework. Neural populations have been shown to optimize their representations for efficient information transmission [linsker1988Selforganization], and Hebbian-like plasticity rules can emerge from that optimization process [linsker1988Selforganization, toyoizumi2005Generalized, toyoizumi2005Spiketiming, toyoizumi2006Fisher]. While STP has been shown to enhance information processing in various contexts [rotman2011Shortterm, regehr2012ShortTerm], the impact of associatively learning STP parameters on neural encoding remains unexplored. In this article, we focus on the functional significance of associative STP from the perspective of efficient neural representations. In particular, we use Fisher information as a local and tractable measure of encoding efficiency [dayan2001Theoretical, brunel1998Mutual].
To this end, we extend the Fisher information optimization framework [seung1993Simple, toyoizumi2006Fisher] from static synaptic weights to dynamic synapses characterized by activity-dependent vesicle release. Specifically, we adopt the Tsodyks-Markram model [tsodyks1997Neural, tsodyks1998Neural] of short-term synaptic dynamics and derive learning rules for both baseline synaptic weight and release probability. We find that this optimization leads to temporal asymmetry: anti-causal connections (where presynaptic neurons lag behind postsynaptic neurons) are preferentially strengthened, contrasting sharply with classical STDP. When stimulus drive varies on timescales slower than the EPSP time constant, the resulting networks emphasize stimulus offsets through “ramp-up” representations and naturally support “reversal” connections of experienced sequences. Furthermore, the extent of temporal asymmetry depends on constraints on release probability, potentially explaining state-dependent differences in replay directionality observed during wakefulness versus sleep. Our framework provides a principled understanding of how associative short-term plasticity shapes neural representations and suggests its previously unappreciated role in rapid learning and memory consolidation.
@@ 2.1 Neuron Model
Following the setting of [seung1993Simple, toyoizumi2006Fisher], we consider a network of stochastically firing spiking neurons.
@@ 2.2 Synapses with Short-Term Dynamics
Synaptic strengths $w_{ij}(t)$ follow Tsodyks-Markram short-term plasticity dynamics [tsodyks1997Neural, tsodyks1998Neural]. In this model, the synaptic efficacy is expressed as the product of a facilitation factor $u_{ij}(t)$ and a depression factor $d_{ij}(t)$:
@@ 2.3 Fisher Information
As calculated in [toyoizumi2006Fisher], the Fisher information, which quantifies how accurately a downstream decoder can estimate $\theta$ from the population activity, is given by:
@@
While [toyoizumi2006Fisher] derived learning rules for time-invariant synaptic weights that maximize Fisher information, our goal is to extend this framework to synapses with short-term dynamics. Specifically, we aim to derive learning rules for both the baseline synaptic strength $w_{ij}^{0}$ and the release probability $U_{ij}$ that maximize the Fisher information in networks with dynamic synapses. Although the initial postsynaptic response scales with the overall product of $w_{ij}^{0}U_{ij}$, these parameters play distinct roles: $U_{ij}$ additionally governs how synaptic efficacy evolves during presynaptic spike trains through STP dynamics, justifying their separate optimization.
@@ 3.1.1 Exact gradient for arbitrary recurrent coupling
Equations (8)–(13) provide an exact representation of $\partial J/\partial Z_{ij}$ for finite recurrent coupling. A complete derivation is given in Appendix A.
@@ 3.1.2 Computability from network simulations
Importantly, the right-hand side of (8) is a function of quantities available along a single simulated trajectory $X(T)$: $\rho_{i}(t)$, $g_{i}^{\prime}(t)/g_{i}(t)$, and the eligibility trace $e_{ij}^{Z}(t)$. The trace $e_{ij}^{Z}(t)$ can be updated online during simulation by differentiating the membrane-potential dynamics with respect to $Z_{ij}$. For the exponential EPSP kernel $\epsilon(t)=e^{-t/\tau_{m}}\Theta(t)$, the eligibility obeys a simple event-driven recursion (continuous decay between presynaptic spikes, with jumps at $t=t_{j}^{f}$ proportional to $\partial_{Z_{ij}}w_{ij}(t_{j}^{f})$), while the latter is obtained from the STP state variables and their parameter sensitivities. We provide explicit update equations for STD synapses in Appendix B. Thus, an unbiased Monte Carlo estimator of $\partial J/\partial Z_{ij}$ is obtained by averaging (8) over repeated simulations (or, in stationary/periodic regimes, by time-averaging after transients).
@@ 3.1.3 Weak-coupling reduction
While (8) is exact, it typically requires Monte Carlo averaging. Following [toyoizumi2006Fisher], we obtain a compact analytic form by assuming weak synaptic coupling, $w_{ij}^{0}\ll 1$ and consequently $w_{ij}(t)\ll 1$, and expanding around the baseline state $w_{ij}^{0}\equiv 0$. At $w=0$, the conditional intensity reduces to a deterministic function
@@
and $\mathcal{J}[X]$ becomes non-random; consequently, the score term in (8) vanishes at leading order. Using the compensation formula for inhomogeneous Poisson spiking at $w=0$ (Appendix C), the gradient reduces to
$\frac{\partial J}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\,\nu_{i}^{0}(t)\eta_{i}(t)\int_{0}^{t}dt^{\prime}\,\epsilon(t-t^{\prime})\,\frac{\partial\langle w_{ij}(t^{\prime})\rangle_{X(t^{\prime})}}{\partial Z_{ij}}\,\nu_{j}^{0}(t^{\prime})\qquad(Z_{ij}=w_{ij}^{0},\,U_{ij}),$ (15)
@@
Under weak coupling, these functions obey closed dynamics induced by the STD model (Appendix C.3):
@@
We further quantified this temporal shift by analyzing the linear response to weak sinusoidal modulation. Figure 2 illustrates the frequency response of the effective presynaptic term $C^{Z}(t)$, summarized by its complex gain $H_{C_{\mathrm{pre}}}^{Z}(\omega)$, whose magnitude and phase give the amplitude gain and phase shift relative to the input (see Appendix D.2 for details).
@@
Crucially, the phase lead for $U$ is consistently larger than for $w_{0}$ across relevant frequencies. Also, the frequencies at which these responses peak ($\omega_{*}$) differ between parameters, implying that $w_{0}$ and $U$ are tuned to extract different temporal features of the input statistics. A more nuanced characterization of how these gains and phase relationships depend on frequency, baseline rate, and STP parameters is presented in Appendix D.2.
@@
A natural implication of our Fisher-information optimization analysis in the above sections is the emergence of anti-causal feedback connectivity patterns that could support reverse replay–like activity. At the level of individual synapses, the postsynaptic factor $\nu_{i}^{0}(t)\eta_{i}(t)$ represents the local Fisher-information contribution and typically scales with the postsynaptic firing rate. Consequently, the learning rule reinforces connections where the onset of presynaptic activity (high $C^{Z}_{ij}$) predicts high postsynaptic information (high $\nu_{i}\eta_{i}(t)$). Because the presynaptic term is effectively phase-advanced (peaking before the rate maximum), this learning rule naturally favors anti-causal associations, i.e., connections from neurons that fire later to neurons that fired earlier in the stimulus-evoked sequence. This tendency is consistent for both the release probability $U_{ij}$ and the baseline weight $w^{0}_{ij}$, but the stronger phase lead in $U_{ij}$ suggests that the optimization of synaptic reliability is particularly driven by rapid temporal changes in the inputs (approximately the temporal derivative), potentially enabling the network to encode dynamic stimulus features more efficiently than static firing rates. At the network level, such anti-causal feedback projections imply that neurons with later preferred positions or phases tend to drive neurons with earlier preferences. This backward connectivity motif has been shown to support reverse replay in hippocampal place-cell models [haga2018Recurrent].
@@ 3.2.2 Optimization of synaptic parameters
Under the same weak-coupling approximation, we can display the optimal profiles explicitly in close analogy to [toyoizumi2006Fisher]. For $w_{0}(\zeta)$, we impose a zero-mean (balance) constraint and a fixed variance (synaptic cost), which yields an optimal profile that is proportional to the gradient shape:
@@ 3.2.3 Constraint-dependent modulation of temporal asymmetry
3.2.4 Changes in Neural Representation
To assess how the learning rules reshape neural representations, we examine the recurrent synaptic input and the resulting firing-rate profile (Figure 7). After learning, the synaptic input current is concentrated in periods of postsynaptic activity and remains close to zero when the postsynaptic neuron is inactive. Within the active window, the input becomes biased toward the latter half of the bump. Correspondingly, the firing-rate bump shifts its peak to later times and exhibits asymmetry between onset and offset. These changes indicate that learning redistributes synaptic gain within the bump, selectively emphasizing later phases of the response instead of uniformly amplifying activity.
We next compare different optimization conditions. When both $U$ and $w_{0}$ are plastic, the representation exhibits the clearest temporal asymmetry and the largest contrast between active and inactive periods. If $U$ is held fixed and only $w_{0}$ is optimized, the synaptic input still grows during bump activity, but the bias toward the latter half and the resulting temporal asymmetry are substantially weakened. Conversely, when $U$ remains plastic but $w_{0}$ is fixed, the bump becomes more asymmetric in time, but the overall modulation amplitude is reduced. Finally, in networks without STP dynamics, synaptic input lacks temporal asymmetry and instead develops negative components during quiescent periods. This inhibitory input outside the postsynaptic activity window, however, barely affects the bump profile, indicating that recurrent resources are spent without effective modulation of the neural activity. In contrast, STP-based models confine synaptic input to task-relevant epochs, thereby achieving more efficient representations for a given synaptic cost.
Figure 7: Changes in neural representation induced by synaptic optimization. Upper panel: recurrent synaptic input $\int_{-\pi}^{\pi}d\zeta\int dt^{\prime}\,\epsilon(t-t^{\prime})w(\zeta,t)\,\nu^{0}(-\zeta)$ to a representative neuron. Lower panel: corresponding firing-rate bump $\nu(t)$. Blue: joint optimization of $U$ and $w_{0}$ (plastic STP and baseline weight). Purple: optimization of $w_{0}$ with fixed $U$. Light blue: optimization of $U$ with fixed $w_{0}$. Green: optimization of a static synapse $w$ without STP dynamics. Black dash (in lower panel): original firing rate $\nu_{0}(t)$.
3.2.5 Fisher Information Enhancement
We next quantify how these representation changes affect the Fisher information. For small synaptic modifications, the change in Fisher information can be approximated as
$\Delta J\approx\int_{0}^{T}dt\int_{-\pi}^{\pi}dz\,h^{\prime}(z,t)\left(\frac{g^{\prime 0}(z,t)}{g^{0}(z,t)}\right)^{2}\Delta\nu(z,t),$ (25)
where $\Delta\nu(z,t)$ denotes the learning-induced change in firing rate.
Because an overall increase in firing rate trivially elevates Fisher information in the linear-response regime, we compare different optimization schemes under a fixed magnitude of membrane-potential perturbation. Specifically, we constrain the total deviation
$\mathcal{V}=\int_{0}^{T}dt\int_{-\pi}^{\pi}dz\,[\Delta u(z,t)]^{2},$ (26)
where $\Delta u(z,t):=u(z,t)-h(z,t)$ is the change in membrane potential induced by recurrent input. Figure 8 shows the resulting Fisher-information gain $\Delta J/J^{0}$ for each condition, normalized by $\mathcal{V}$.
Among synapses with and without STP, Fisher information is maximally enhanced when both $U$ and $w_{0}$ are plastic. In this case, associative learning jointly shapes presynaptic release dynamics and postsynaptic gain, confining synaptic input to the bump while suppressing it during quiescent periods. This selective gating allows the network to approach an optimal trade-off between information enhancement and deviation from the original activity pattern, effectively maximizing information gain per unit of metabolic or homeostatic cost.
Figure 8: Fisher information enhancement under different optimization schemes. Relative change in Fisher information $\Delta J/J^{0}$ across optimization conditions, normalized by the total membrane-potential fluctuation $\mathcal{V}$. Synapses with STP achieve the largest information gain when both $U$ and $w_{0}$ are plastic.

### v2

@@
Abstract
Short-term synaptic plasticity (STP) is often regarded as a presynaptic filter of spikes, independent of postsynaptic activity. Recent experiments, however, indicate an associative STP that depends on pre- and postsynaptic coactivation. We develop a normative, information-theoretic theory of associative STP. Extending Fisher-information-based learning to Tsodyks–Markram synapses, we derive learning rules for baseline weight and release probability that maximize stimulus information under resource constraints. The rules split into a postsynaptic term tracking local firing and a presynaptic, phase-advanced term that selectively detects stimulus onset. For slowly varying inputs, this onset sensitivity favors anti-causal connectivity and enhances response offset during drive and reverse replay after drive removal in recurrent circuits. Linear-response analysis shows that STP yields frequency-dependent phase selectivity and that release-probability constraints tune temporal asymmetry. These results identify release-probability plasticity as a principled substrate for rapidly reconfigurable temporal coding.
@@ 1 Introduction
Learning through experience is mediated by synaptic plasticity, the activity-dependent modification of connection strengths between neurons. Classical long-term potentiation (LTP) has been extensively studied as the primary mechanism for associative learning, where correlated pre- and postsynaptic activity leads to persistent strengthening of synapses through increased postsynaptic receptor expression (?, ?). This Hebbian plasticity enables the formation of neural assemblies that encode relationships between concepts. However, LTP typically requires tens of minutes to hours for full expression and once expressed, can persist from hours to days or longer, which cannot fully account for flexible behavioral adaptation occurring on faster timescales.
@@
To address this question, we adopt an information-theoretic framework. Neural populations have been shown to optimize their representations for efficient information transmission (?), and Hebbian-like plasticity rules can emerge from that optimization process (?, ?, ?, ?). While STP has been shown to enhance information processing in various contexts (?, ?), the impact of associatively learning STP parameters on neural encoding remains unexplored. In this article, we focus on the functional significance of associative STP from the perspective of efficient neural representations. In particular, we use Fisher information as a local and tractable measure of encoding efficiency (?, ?).
To this end, we extend the Fisher information optimization framework (?, ?) from static synaptic weights to dynamic synapses characterized by activity-dependent vesicle release. Specifically, we adopt the Tsodyks-Markram model (?, ?) of short-term synaptic dynamics and derive learning rules for both baseline synaptic weight and release probability. We find that this optimization leads to temporal asymmetry: anti-causal connections (where presynaptic neurons lag behind postsynaptic neurons) are preferentially strengthened, contrasting sharply with classical STDP. When stimulus drive varies on timescales slower than the EPSP time constant, the resulting networks emphasize stimulus offsets through “ramp-up” representations and naturally support “reversal” connections of experienced sequences. Furthermore, the extent of temporal asymmetry depends on constraints on release probability, potentially explaining state-dependent differences in replay directionality observed during wakefulness versus sleep. Our framework provides a principled understanding of how associative short-term plasticity shapes neural representations and suggests its previously unappreciated role in rapid learning and memory consolidation.
@@ 2.1 Neuron Model
Following the setting of (?, ?), we consider a network of stochastically firing spiking neurons.
@@ 2.2 Synapses with Short-Term Dynamics
Synaptic strengths $w_{ij}(t)$ follow Tsodyks-Markram short-term plasticity dynamics (?, ?). In this model, the synaptic efficacy is expressed as the product of a facilitation factor $u_{ij}(t)$ and a depression factor $d_{ij}(t)$:
@@ 2.3 Fisher Information
As calculated in (?), the Fisher information, which quantifies how accurately a downstream decoder can estimate $\theta$ from the population activity, is given by:
@@
While (?) derived learning rules for time-invariant synaptic weights that maximize Fisher information, our goal is to extend this framework to synapses with short-term dynamics. Specifically, we aim to derive learning rules for both the baseline synaptic strength $w_{ij}^{0}$ and the release probability $U_{ij}$ that maximize the Fisher information in networks with dynamic synapses. Although the initial postsynaptic response scales with the overall product of $w_{ij}^{0}U_{ij}$, these parameters play distinct roles: $U_{ij}$ additionally governs how synaptic efficacy evolves during presynaptic spike trains through STP dynamics, justifying their separate optimization.
@@ 3.1.1 Exact gradient for arbitrary recurrent coupling
Equations (8)–(13) provide an exact representation of $\partial J/\partial Z_{ij}$ for finite recurrent coupling. A complete derivation is given in Appendix 5.1.
@@ 3.1.2 Computability from network simulations
Importantly, the right-hand side of (8) is a function of quantities available along a single simulated trajectory $X(T)$: $\rho_{i}(t)$, $g_{i}^{\prime}(t)/g_{i}(t)$, and the eligibility trace $e_{ij}^{Z}(t)$. The trace $e_{ij}^{Z}(t)$ can be updated online during simulation by differentiating the membrane-potential dynamics with respect to $Z_{ij}$. For the exponential EPSP kernel $\epsilon(t)=e^{-t/\tau_{m}}\Theta(t)$, the eligibility obeys a simple event-driven recursion (continuous decay between presynaptic spikes, with jumps at $t=t_{j}^{f}$ proportional to $\partial_{Z_{ij}}w_{ij}(t_{j}^{f})$), while the latter is obtained from the STP state variables and their parameter sensitivities. We provide explicit update equations for STD synapses in Appendix 5.2. Thus, an unbiased Monte Carlo estimator of $\partial J/\partial Z_{ij}$ is obtained by averaging (8) over repeated simulations (or, in stationary/periodic regimes, by time-averaging after transients).
@@ 3.1.3 Weak-coupling reduction
While (8) is exact, it typically requires Monte Carlo averaging. Following (?), we obtain a compact analytic form by assuming weak synaptic coupling, $w_{ij}^{0}\ll 1$ and consequently $w_{ij}(t)\ll 1$, and expanding around the baseline state $w_{ij}^{0}\equiv 0$. At $w=0$, the conditional intensity reduces to a deterministic function
@@
and $\mathcal{J}[X]$ becomes non-random; consequently, the score term in (8) vanishes at leading order. Using the compensation formula for inhomogeneous Poisson spiking at $w=0$ (Appendix 5.3), the gradient reduces to
$\frac{\partial J}{\partial Z_{ij}}=\int_{0}^{T}dt\,\nu_{i}^{0}(t)\eta_{i}(t)\int_{0}^{t}dt^{\prime}\,\epsilon(t-t^{\prime})\,\\ \frac{\partial\langle w_{ij}(t^{\prime})\rangle_{X(t^{\prime})}}{\partial Z_{ij}}\,\nu_{j}^{0}(t^{\prime})\qquad(Z_{ij}=w_{ij}^{0},\,U_{ij}),$ (15)
@@
Under weak coupling, these functions obey closed dynamics induced by the STD model (Appendix 5.3.3):
@@
We further quantified this temporal shift by analyzing the linear response to weak sinusoidal modulation. Figure 2 illustrates the frequency response of the effective presynaptic term $C^{Z}(t)$, summarized by its complex gain $H_{C_{\mathrm{pre}}}^{Z}(\omega)$, whose magnitude and phase give the amplitude gain and phase shift relative to the input (see Appendix 5.4.2 for details).
@@
Crucially, the phase lead for $U$ is consistently larger than for $w_{0}$ across relevant frequencies. Also, the frequencies at which these responses peak ($\omega_{*}$) differ between parameters, implying that $w_{0}$ and $U$ are tuned to extract different temporal features of the input statistics. A more nuanced characterization of how these gains and phase relationships depend on frequency, baseline rate, and STP parameters is presented in Appendix 5.4.2.
@@
A natural implication of our Fisher-information optimization analysis in the above sections is the emergence of anti-causal feedback connectivity patterns that could support reverse replay–like activity. At the level of individual synapses, the postsynaptic factor $\nu_{i}^{0}(t)\eta_{i}(t)$ represents the local Fisher-information contribution and typically scales with the postsynaptic firing rate. Consequently, the learning rule reinforces connections where the onset of presynaptic activity (high $C^{Z}_{ij}$) predicts high postsynaptic information (high $\nu_{i}\eta_{i}(t)$). Because the presynaptic term is effectively phase-advanced (peaking before the rate maximum), this learning rule naturally favors anti-causal associations, i.e., connections from neurons that fire later to neurons that fired earlier in the stimulus-evoked sequence. This tendency is consistent for both the release probability $U_{ij}$ and the baseline weight $w^{0}_{ij}$, but the stronger phase lead in $U_{ij}$ suggests that the optimization of synaptic reliability is particularly driven by rapid temporal changes in the inputs (approximately the temporal derivative), potentially enabling the network to encode dynamic stimulus features more efficiently than static firing rates. At the network level, such anti-causal feedback projections imply that neurons with later preferred positions or phases tend to drive neurons with earlier preferences. This backward connectivity motif has been shown to support reverse replay in hippocampal place-cell models (?).
@@ 3.2.2 Optimization of synaptic parameters
Under the same weak-coupling approximation, we can display the optimal profiles explicitly in close analogy to (?). For $w_{0}(\zeta)$, we impose a zero-mean (balance) constraint and a fixed variance (synaptic cost), which yields an optimal profile that is proportional to the gradient shape:
@@ 3.2.3 Constraint-dependent modulation of temporal asymmetry
3.2.4 Associative STP induces ramping representations
The weak-coupling analysis above provides intuition for how associative STP biases learning toward anti-causal associations. We next test whether these predictions persist in a fully recurrent spiking network by optimizing Fisher information using the exact, simulation-based gradient estimator (Secs. 3.1.1 and 3.1.2) in the traveling-wave setting (Sec. 3.2).
We compare three synaptic models: (i) associative STP, in which both the baseline weight profile $w_{0}(\zeta)$ and the release-probability profile $U(\zeta)$ are optimized; (ii) non-associative STP, in which synapses exhibit Tsodyks–Markram dynamics but $U(\zeta)$ is held fixed and only $w_{0}(\zeta)$ is learned; and (iii) static synapses without STP dynamics.
After learning, networks with associative STP exhibit a pronounced ramping of stimulus-evoked activity: within each stimulus epoch, firing rates increase toward the latter part of the traveling bump, effectively emphasizing stimulus offsets (Figure 7A). In contrast, non-associative STP and static synapses do not develop robust ramping; instead, learning primarily redistributes activity away from the instantaneous input peak toward its flanks (Figure 7B,C).
To relate these activity changes to circuit interactions, we computed pairwise cross-correlograms during both stimulus-evoked and spontaneous activity as a proxy for effective coupling (Figure 8). Associative STP strengthens interactions in the direction opposite to the experienced propagation (anti-causal), consistent with the temporal bias predicted by our analytic learning rule (Figure 8A,D). Non-associative STP produces a qualitatively similar but weaker and less specific pattern, with a narrower spread in both time and space (Figure 8B,E), whereas static synapses show no clear emergence of structured correlations (Figure 8C,F). Together, these results suggest that plasticity of release probability is critical for converting Fisher-information optimization into ramping representations and robust sequence-specific effective connectivity, motivating our analysis of reverse replay next.
3.2.5 Associative STP supports reverse replay during spontaneous activity
The results above indicate that learning with associative STP strengthens effective coupling in the direction opposite to the experienced stimulus propagation. Such “backward” connectivity motifs have been proposed as a circuit mechanism for hippocampal reverse replay (?). We therefore asked whether Fisher-information optimization in our model gives rise to reverse replay when external drive is removed.
After learning, we simulated the network dynamics in the absence of stimulus drive and examined the resulting spontaneous activity patterns. Networks with STP dynamics generated spontaneous sequential events that propagated opposite to the original stimulus direction, consistent with reverse replay (Figure 9A,B). Reverse replay was observed under both associative STP and non-associative STP, but it was markedly more temporally compressed (faster propagation) and more spatially selective (sharper contrast) when release probability was also optimized. In contrast, networks with static synapses did not exhibit reliable replay-like sequences (Figure 9C).